Whirly Corporations contribution format income statement for the most recent month is shown below: Sales (7,200 units) Variaplease answer all parts. Thanks!
Answer #1

Revised net operating income

= (7200+100)*15 - 54,700

= 54,800

Revised net operating income

= (7200-100)*15 - 54,700

= 51,800

Revised net operating income

= (6200*15) - 54,700

= 38,300